SOLVE THE EQUATION AND CHECK THE SOLUTION


X + 8 = 2 ( X - 2) - X

Answer by Mathtut(3670)


X + 8 = 2 ( X - 2) - X
:
first distribute the right side
:
x+8=2x-4-x
:
x+8=x-4--->if you subtract x from both sides we have 8=-4
:
this statement is obviously not true therefore there is NO SOLUTION